OpenStudy (anonymous):

The elevation of city A is 80 feet below sea level; the elevation of city C is 16 feet below sea level. The elevation of city D is 1 over 2 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. (5 points) Part B: Solve the expression and explain why the answer is negative or positive.

OpenStudy (hybrik):

D = 1/2(C)

OpenStudy (hybrik):

Part B: D=(1/2*16), D = 8

OpenStudy (anonymous):

@hybrik Please do not give direct answers that is against the OS COC.

OpenStudy (anonymous):

yea i just wanted help

OpenStudy (anonymous):

So @beckyg111 , what do you have so far?

OpenStudy (anonymous):

wdym?

OpenStudy (anonymous):

Like wat do you have written so far?

OpenStudy (hybrik):

"exact"

OpenStudy (anonymous):

??

OpenStudy (hybrik):

"direct answers"

OpenStudy (hybrik):

I was going to explain..

OpenStudy (anonymous):

lol no

OpenStudy (hybrik):

Okay so part A, C is 16, using 1/2 as said "one over two" you have to multiply 1/2 * 16

OpenStudy (anonymous):

You did give a direct answer. You do not give the answer and explain. You explain and let them find the answer it makes them feel more accomplished. Please do not listen to @hybrik

OpenStudy (hybrik):

What a great way to get medals.

OpenStudy (anonymous):

amen!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So anyways, what is the last sentence saying?

OpenStudy (anonymous):

ok let me read the question again

OpenStudy (anonymous):

k, i submitted my test got a 100 thanks for the help @barbieprincess :)

OpenStudy (anonymous):

Smh some people do not know how to follow rules.

OpenStudy (anonymous):

i did my work though i didnt cheat btw

OpenStudy (anonymous):

I am trying to explain this @beckyg111 but if you want to listen to answers to get your A's you will never pass the SAT or ACT

OpenStudy (anonymous):

js

OpenStudy (anonymous):

no no, i'm not cheating lol. I know this will only hurt me if i do

OpenStudy (anonymous):

i did do my work.

OpenStudy (anonymous):

But you did not understand it and I never explained how did you get an A?

OpenStudy (anonymous):

well i have other questions besides the one above. My dad just came in and explained it to me. Look, sometimes i only look at the question once and give up.I just read it like 4 times.

OpenStudy (anonymous):

then i had that "ohhhh" moment

OpenStudy (anonymous):

Me too. Lol okay good luck.